If sin(x+30)=1/2 (0≤x≤360), then x is equal to?
My attempt-
sin(x+30)=1/2
sin(angle)=1/2
shift sin 1/2=30
x+30=30
x=0
so x=0, 180, 360
this is a multiple choice question and the answers are-(A) 0,120 OR 360. (B) 0 OR 120 (C) 30 OR 150 (D) 120 OR 360
Thanks for any help in advance

Answer by stanbon(57239)

If sin(x+30)=1/2 (0≤x≤360), then x is equal to?
----
If sin(x+30) = 1/2, x+30 must be 30, or 150, or 390
----
If x+30 = 30, x = 0
If x+30 = 150, x = 120
If x+30 = 390, x = 360
Ans: A
---
Cheers,
Stan H.